OpenStudy (anonymous):

MEDAL!! what function rule is represented by the table? (table will be posted momentarily)

OpenStudy (anonymous):



OpenStudy (anonymous):

@dan815 @dtan5457 @jim_thompson5910 @jdoe0001

OpenStudy (anonymous):

x is the position of y

OpenStudy (anonymous):

the choices are= a) y=x+3 b) y=x+6 c) y=2x+3 or... d) y=4x+1

OpenStudy (radar):

Just check which ones satisfies the table......Or think about it and solve. You could start by answering jim_thompson5910 question and then determine how much y is increasing>

OpenStudy (jdoe0001):

\(\bf \begin{array}{rllllll} x&&0&{\color{red}{ 1}}&2&{\color{red}{ 3}}\\ y&&3&{\color{blue}{ 5 }}&7&{\color{blue}{ 9}} \end{array} \qquad \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 1}}\quad ,&{\color{blue}{ 5}})\quad % (c,d) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

get the slope, plug the values, solve for "y" that'd give you the equation as jim_thompson5910 said, the change will be the slope, and that'd give you a start

OpenStudy (anonymous):

I know y is increasing by +2 each time but I don't know which rule it would be from the choices #confused

OpenStudy (radar):

Don't forget slope is change in y divided by change in x. Also slope will be the coefficient of x in the function equation.

OpenStudy (anonymous):

wait is the answer d? cause I tried a different way and it makes more sense for this one to be the answer

OpenStudy (radar):

That was a guess.....

OpenStudy (radar):

Ummm let me see d?? y = 4x + 1 then when x = 0, y would = 1 but table says y = 3 BBBZZTT wrong answer.

OpenStudy (anonymous):

lol ok so, when x =0 y= 2(0) +3=3 and the table says 3 so is it c?

OpenStudy (radar):

Well that looks better but to be sure check it with the next value of x......try x=1.

OpenStudy (radar):

does c work with x=1?

OpenStudy (anonymous):

ok, when x=1 y=2(1)+3= 5 and the table says 5 so it is c! right?

OpenStudy (radar):

I am convinced, but the cange in y is 2 when the change in x is 1 so the slope is 2/1 or 2, that means y= 2x + b, and the intercept b occurs when x is 0. I would say all the evidence is pointing at C,

OpenStudy (radar):

Are you convinced it is c.

OpenStudy (anonymous):

ok, and yes im convinced it is c, Thank You :D

OpenStudy (radar):

My pizza is ready, gotta run. good luck with your studies and with your kitten, my main pet is the Boston Terrier on my lap in the photo.